OpenStudy (anonymous):

Please Help solve for x: 4/x-3 + 5/x+3=6/x^2-9

OpenStudy (anonymous):

(x^2-9 ) is the difference of two squares therefore it can also be expressed as (x-3) (x+3) bring the left hand side of the equation under the same denominator

OpenStudy (anonymous):

you should get 9x-3/(x-3)(x+3) = 6/ (x-3)(x+3)

OpenStudy (anonymous):

this is the answer right

OpenStudy (anonymous):

what i just posted no x=1

OpenStudy (anonymous):

so x=1 is the answer

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

multiply all what by (x-3)(x+3)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

ok but thats not the answer right

jhonyy9 (jhonyy9):

4 5 6 ----- + ------- = -------- x -3 x +3 x^2 -9 first you need getting the common denominator so what is x^2 -9 what you can rewrite it like (x-3)(x+3) exactly how have wrote above @hammytams so than now you need multiplie the numerator on the first fraction by (x+3) and the numerator on the second fraction by (x+3) so in this way you will get 4(x+3) 5(x-3) 6 ---------- + ---------- = ---------- (x-3)(x+3) (x-3)(x+3) (x-3)(x+3) so than 4(x+3) + 5(x-3) = 6 4x+12 +5x -15 =6 9x -3 = 6 9x = 9 x= 9/9 x=1